Stickybuns:

Select the correct answer from each drop-down menu. Consider the function f(x) = 3x + 1 and the graph of the function g(x) shown below.

Stickybuns:

1 attachment
Stickybuns:

The graph g(x) is the graph of f(x) translated___ units____ , and g(x) = ____

Stickybuns:

first fill in is 5 3 and 2

Stickybuns:

2nd is left right up and down

Angle:



Angle:

We should look specifically at where each graph intersects the y-axis

Angle:

f(x) crosses the y-axis at 1 g(x) crosses the y-axis at -5

Stickybuns:

3rd is f(x)-5 f(x)+2 f(x-2) f(x) +3 f(x-3) f(x+5)

Angle:

ok, let's consider the first fill-in f(x) crosses the y-axis at 1 g(x) crosses the y-axis at -5 how much is that distance?

Stickybuns:

5? right cause -5 -4 -3 -2 -1 0

Stickybuns:

not including the -5 its 5

Angle:

huh... this is odd... the answer is not one of the options? o-o

Angle:

:checks again: yeah, no,... f(x) is definitely 6 up from g(x)

Angle:

-5, -4, -3, -2, -1, 0, 1 is a difference of 6

Stickybuns:

hmmm do you think they could mean not including -5?

Angle:



Stickybuns:

hmm then we have a problem

Angle:

indeed... could you check that f(x) = 3x + 1 or f(x) = 3x - 1 ? or the g(x) picture to see if that's the right one?

Stickybuns:

f(x) = 3x + 1

Angle:

whelp...

Angle:

I guess if you need to put down an answer for now, 5 is the closest... but you should probably ask your teacher about that

Stickybuns:

Okay will do

Stickybuns:

the 2nd fill in would be up right?

Angle:

well f(x) is higher than g(x) and the question is asking g(x) in relation to f(x) so g(x) is under f(X)

Stickybuns:

under = down so down ?

Angle:

correct ^_^

Stickybuns:

now for the hard looking part

Angle:

f(x)-5 = g(x) is 5 units below f(x) f(x)+2 = g(x) is 2 units above f(x) f(x-2) = g(x) is 2 units to the right of f(x) f(x) +3 = g(x) is 3 units above f(x) f(x-3) = g(x) is 3 units to the right of f(x) f(x+5) = g(x) is 5 units to the left of f(x)

Stickybuns:

mmm so hard XD a? for once again the below?

Angle:

lol correct

Stickybuns:

some how that was all wrong XD

Stickybuns:

Consider the function d(x) with real numbers h > 0 and k > 0. The function g(x) can be seen as a transformation of the function f(x). First, find the y-intercept of the function f(x) = 3x + 1. The function is written in slope-intercept form, so the y-intercept of f(x) is (0, 1). Find the point where the function g(x) also has an output value of 1. This point is (2, 1) and is 2 units to the right of the y-intercept of f(x). A horizontal shift to the right can be written as d(x - h). Therefore, the graph g(x) is the graph of f(x) translated 2 units right, and g(x) = f(x - 2

Stickybuns:

1 attachment
Stickybuns:

https://media.studyisland.com/cgi-bin/mimetex.cgi?\begin{array}{rclC30C30C30C30}%20d(x)%20\%20+%20\%20k%20\%20&\rightarrow&%20\%20\text{shifts}%20\%20d(x)%20\%20\text{up}%20\%20k%20\%20\text{units}\\%20d(x)%20\%20-%20\%20k%20\%20&\rightarrow&%20\%20\text{shifts}%20\%20d(x)%20\%20\text{down}%20\%20k%20\%20\text{units}\\%20d(x%20\%20+%20\%20h)%20\%20&\rightarrow&%20\%20\text{shifts}%20\%20d(x)%20\%20\text{left}%20\%20h%20\%20\text{units}\\%20d(x%20\%20-%20\%20h)%20\%20&\rightarrow&%20\%20\text{shifts}%20\%20d(x)%20\%20\text{right}%20\%20h%20\%20\text{units}\\%20\end{array}

Stickybuns:

nurp okay XD